wat will be d integration of

tan(x-k).tan(x+k).tan2xdx.......???

Grade:12

1 Answers

Rahul askIITiansExpert.IITR
13 years ago

very simple question

see:

tan(2x)=tan((x-k)+(x+k))=tan(x-k) + tan(x+k)/1 + tan(x-k)tan(x+k)

take the denominator to right side and you will get the expression in the question. Integrate it.

 

Got it...If yes please approve..